OpenStudy (anonymous):

Can someone please help me with this problem? (8x/x^2-4)/(4x^2/x^2+4x+4)

OpenStudy (anonymous):

\[\large \frac{8x}{x^2-4}/\frac{4x^2}{x^2+4x+4}=\frac{8x}{x^2-4}\times\frac{x^2+4x+4}{4x^2} \]

OpenStudy (anonymous):

try factorizing

OpenStudy (anonymous):

I already tried, I have no idea what to do.

OpenStudy (anonymous):

8*(x+2)^2/4x*[(x^2) -4]

OpenStudy (pfenn1):

Hint: \[x^2-c^2 = (x+c)(x-c)\]

OpenStudy (anonymous):

\[8*(x+2)^2/4x^3-16x\]

OpenStudy (pfenn1):

\[x^2+4x+4=(x+a)(x+b)\]where a and b are integers where (a + b) = 4 and ab = 4.

OpenStudy (anonymous):

what's the answer?

OpenStudy (anonymous):

try finding it yourself.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\[x^2+4x+4=(x+2)^2\]

OpenStudy (anonymous):

\[x^2-4=(x+2)(x-2)\]

OpenStudy (anonymous):

\[8*x*(x+2)^2/(x^2-4)*4x^2=2(x+2)^2/x*(x^2-4)\]

OpenStudy (anonymous):

\[2*(x+2)^2/x*(x+2)(x-2)=2*(x+2)/x(x-2)\]

OpenStudy (anonymous):

understand*

OpenStudy (anonymous):

`yep